OpenStudy (anonymous):

Use the irregular polygon below. Show all work MEDAL!! FAN!!!! HELP!!!! a. Find the perimeter. b. Find the area

OpenStudy (anonymous):

all lines are straight

OpenStudy (anonymous):



OpenStudy (michele_laino):

perimeter is the sum of the lengths of the sides, so: perimeter = 8+8+4+2+(8-2)+(8-4)=...?

OpenStudy (anonymous):

50??

OpenStudy (anonymous):

or is it 32?

OpenStudy (michele_laino):

yes! it is 32 mm

OpenStudy (anonymous):

ok thanks!!

OpenStudy (anonymous):

what about the area?

OpenStudy (michele_laino):

In order to compute the are, we have to compute this: Area= 8*4+ (8-4)*2=...? since our polygon can be viewed as two rectangles

OpenStudy (michele_laino):

oops,.. in order to compute the area...

OpenStudy (anonymous):

32.2

OpenStudy (michele_laino):

Hint: 8*4 + (8-2) *2= 32+ 6*2=32+12=...?

OpenStudy (anonymous):

hold on brb

OpenStudy (anonymous):

@Michele_Laino 44?

OpenStudy (anonymous):

@kirbykirby @JadeK help

OpenStudy (michele_laino):

that's right! area = 44 mm^2

OpenStudy (anonymous):

ok. thanks so much